In a test of ESP (extrasensory perception), the experimenter looks at cards that are hidden from the subject. Each card contains either a star, a circle, a wavy line, or a square. An experimenter looks at each of 100 cards in turn, and the subject tries to read the experimenter's mind and name the shape on each. What is the probability that the subject gets 35 or more correct if the subject does not have ESP and is just guessing?
Answer #1

Answer:

Given,

sample n = 100

p = 1/4 = 0.25

q = 1 - 0.25 = 0.75

mean = np = 100*0.25 = 25

standard deviation = sqrt(npq) = sqrt(100*0.25*0.75) = 4.33

Here from normal approximation of binomial distribution and continuity correction

P(X >= 34.5) = P((x-u)/s > (34.5 - 25)/4.33)

= P(z > 2.19)

= 0.0142621 [since from z table]

= 0.0143